OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 3 = 2 ?

OpenStudy (anonymous):

\[\log _{2}x-\log _{2}3=2\]

OpenStudy (anonymous):

combine in to one log as a first step, then write in equivalent exponential form

OpenStudy (anonymous):

i'm not sure how to do that.

OpenStudy (anonymous):

use \[\log(x)-\log(y)=\log(\frac{x}{y})\] for the first part

OpenStudy (anonymous):

so step one is to write \[\log_2(\frac{x}{3})=2\]

OpenStudy (anonymous):

you got step 2? equivalent exponential form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good